OpenStudy (anonymous):

Rewrite x2 - 4x + 4 as a square of a linear expression

OpenStudy (amoodarya):

\[x^2+ax =(x +\frac{ a }{ 2 })^2 -(\frac{ a }{ 2 })^2\]